rebeccaxhawaii (rebeccaxhawaii):

What is the slope of the line passing through the points (2, 5) and (0, −4)?

rebeccaxhawaii (rebeccaxhawaii):

Posting on @MissHollywood 's behalf

rebeccaxhawaii (rebeccaxhawaii):



Mehek (mehek14):

Use the slope formula \(\bf{\dfrac{y_2-y_1}{x_2-x_1}}\)

rebeccaxhawaii (rebeccaxhawaii):

Plug in your points.

Mehek (mehek14):

Plug in the values of the ordered pair \(\dfrac{-4-5}{0-2}=?\)

OpenStudy (misshollywood):

I got 9/2

rebeccaxhawaii (rebeccaxhawaii):

Correct